A pitcher's earned-run average E varies directly as the number R of earned runs allowed and inversely as the number I of innings pitched. A certain pitcher had an earned-run average of 2.92, giving up 85 earned runs in 262 innings. How many earned runs would the pitcher have given up having pitched 284 innings with the same average.
Any help with setting this up would be greats. Thanks:)

Answer by stanbon(75887)

A pitcher's earned-run average E varies directly as the number R of earned runs allowed and inversely as the number I of innings pitched. A certain pitcher had an earned-run average of 2.92, giving up 85 earned runs in 262 innings. How many earned runs would the pitcher have given up having pitched 284 innings with the same average.
-------
E = k*R/I
-----
Solve for "k" using "A certain pitcher had an earned-run average of 2.92, giving up 85 earned runs in 262 innings".
2.92 = k*85/262
----
k = 9
----
Equation::
E = 9*R/I
-----
?How many earned runs would the pitcher have given up having pitched 284 innings with the same average?
----
2.92 = 9*R/284
---
R = 92.14 or 92 when rounded down.
Cheers,
Stan H.
--------